Fill in the missing fractions.
 7 / 10 - ▯ = 3 / 10 ???????????

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Given7 / 10 -▯= 3 / 10
▯= 7 / 10–3 / 10
▯= (7–3) / 10
▯= 4 / 10
▯=2 / 5

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
7/10 -  __ = 3/10
Let the missing number be x,
 
7/10 - x = 3/10
 X = 7/10 - 3/10
X = 5/10
 
Hope it helps!
Thank you